Question 2. Consider a variable x that describes the number of under-inflated tire on a four-wheel automobile. The mass function is given by: p(0) 0.4, p(1) p(2) p(3) 0.1; p(4) 0.3 a) Verify that this is a proper mass function. b) Determine the expected value and the variance of this distribution. mean value? please show your work. thanks!  
Answer #1

a)

here as probability of each data point is between 0 and 1 ; and sum of probability over sample space

=P(0)+P(1)+P(2)+P(3)+P(4)=0.4+0.1+0.1+0.1+0.3=1

hence this is valid mass function

b)

x P(x) xP(x) x2P(x)
0 0.4 0.000 0.000
1 0.1 0.100 0.100
2 0.1 0.200 0.400
3 0.1 0.300 0.900
4 0.3 1.200 4.800
total 1.800 6.200
E(x) =μ= ΣxP(x) = 1.8000
E(x2) = Σx2P(x) = 6.2000
Var(x)=σ2 = E(x2)-(E(x))2= 2.960

from above expected value E(X)=1.8

and variance =2.96

c)

as std deviation =sqrt(2.96)=1.7205

values -/+ 1 std deviation from mean =1.8-/+1.7205 =0.08 to 3.52

hence proportion of cars within -/+ 1 std deviation from mean=P(0.08 <X<3.52)=P(1)+P(2)+P(3)=0.1+0.1+0.1=0.3